Solar radiation enters the Earth's atmosphere with a portion being scattered by clouds and aerosols.

Processing, archiving and distributing Earth science data
at the NASA Langley Research Center

ASDC Past Maintenance Events

RSS Feeds - news, announcements, maintenance.

Past Maintenance

Pages

Date(s) Time Event Impact
Wednesday, August 16, 2017 08:00 am - 11:00 am EST

The Data Pool, MISR order and browse tools, TAD, TES and MOPITT Search and Subset Applications will be unavailable from 8am - 11am for scheduled maintenance. Data processing and distribution may be impacted during maintenance. We apologize for the inconvenience.

 

Saturday, July 29, 2017 to Sunday, July 30, 2017 7/29 @ 09:00 am - 7/30 @ 05:00 pm EDT

The ASDC is scheduling a two day outage for the DPO GPFS filesystems and SCF2 file system starting at 9 AM Saturday morning July 29, 2017 through Sunday July 30, 2017.  The Univa Grid queues will be quiesced on AMI and AMI-P the evening of July 28, 2017 at 9PM.  ANGe ingest will be paused beginning at 8am Saturday July 29 and the dropbox will remain online for delivery of data only.  Additionally,  other GPFS mounts (via afp, samba and nfs) may be unavailable as we will need to stop GPFS on the fileservers to perform this maintenance on the DPO and SCF2 filesystems. Data processing and distribution may be impacted during maintenance. We apologize for the inconvenience.

 

Again, the following filesystems will be affected:
ASDC_archive4/5/6/8 (DPO)
CALIPSO, CERES L2, MISR a7 MOPITT Subsetters
OPeNDAP
TAD
SCF2

 

Wednesday, July 19, 2017 07:00 am - 03:00 pm EST

The Data Pool, MISR order and browse tools, TAD, TES and MOPITT Search and Subset Applications, and Reverb will be unavailable from 7am - 3pm for scheduled maintenance. Data processing and distribution may be impacted during maintenance. We apologize for the inconvenience.

 

Wednesday, July 12, 2017 9:00 am - 11:00 am EST

Connectivity to the EOSWeb site will be temporarily unavailable from 9-11 am.

 

Wednesday, March 22, 2017 09:00 am - 09:30 am EDT

Connectivity to the EOSWeb site will be temporarily unavailable from 9-9:30 am.

 

Saturday, March 11, 2017 09:00 am - 05:00 pm EDT

The DPO (archive 4,5,6) GPFS filesystem will be taken offline starting at 9am Saturday March 11, 2017.  The Univa Grid queues will be quiesced on AMI and AMI-P the evening of Friday, March 10, 2017 at 9pm.  ANGe ingest will be paused beginning at 8am Saturday March 11 and the dropbox will remain online for delivery of data only.  This Saturday’s work may potentially impact filesystem (gpfs) access on other hosts. Data processing and distribution may be impacted during the maintenance period. We apologize for any inconvenience this may cause.

 

Wednesday, March 8, 2017 07:00 am - 04:00 pm EST

The Data Pool anonymous ftp server l5eil01.larc.nasa.gov will transition to l5ftl01.larc.nasa.gov on March 8, 2017. Please make any necessary script updates and plan accordingly. MISR NRT data will remain on l5eil01 and can only be accessed through http Earthdata authentication. Issues and concerns should be sent directly to our User Services Team.

 

Access the ASDC DataPool at https://eosweb.larc.nasa.gov/datapool

 

 

Wednesday, March 8, 2017 07:00 am - 04:00 pm EST

Connectivity to the EOSWeb site will be unavailable from 9 - 11 am.

 

The Data Pool, MISR order and browse tools, TAD, TES and MOPITT Search and Subset Applications, and Reverb will be unavailable from 7am - 4pm for scheduled maintenance. Data processing and distribution may be impacted during maintenance. We apologize for the inconvenience.

 

Wednesday, February 8, 2017 08:00 am - 12:00 pm EDT

Connectivity to the EOSWeb site will be unavailable from 9-11 am.

 

The Data Pool, Reverb, MISR order and browse tools, CERES, MOPITT, TAD and TES Search and Subset Applications will be unavailable Wed. from 8am - 12pm for scheduled maintenance. Data processing and distribution may be impacted during the maintenance period. We apologize for any inconvenience this may cause.

 

Saturday, January 28, 2017 09:00 am - 05:00 pm EDT

The ASDC will be performing maintenance on the DPO GPFS filesystems and SCF 2, 5, and 8 starting 9 AM Saturday morning January 28, 2017. ANGe ingest will be paused beginning at 8am Saturday January 28 and the dropbox will remain online for delivery of data only. We anticipate to complete all work in 6 to 8 hours. Data processing and distribution may be impacted during maintenance. We apologize for the inconvenience.

 

Pages